OpenStudy (anonymous):

eliminate the parameter. x = 5 cos t, y = 5 sin t help please!

OpenStudy (amistre64):

which parameter?

OpenStudy (gorv):

@coolkat4 cost=x/5 sint=y/5 sint^2+cost^2=1 (x/5)^2+(y/5)^2=1 x^2+y^2=25

OpenStudy (amistre64):

if it means solve for y in terms of x, solve for t in terms of x

OpenStudy (anonymous):

okay i understand that. now what exactly is eliminating the parameter mean? is that the x^2+y^2=25

OpenStudy (amistre64):

generally, it means creating a: y=______ type setup such that y is a function of x

OpenStudy (anonymous):

so after x^2+y^2=25 where do i go from there?

OpenStudy (gorv):

yes here t is the parameter this the final answer parameters are eliminated

OpenStudy (gorv):

got it????

OpenStudy (anonymous):

kind of..Eliminate the parameter. x = t^2 + 1, y = t^2-1....would the answer be in the same form?

OpenStudy (amistre64):

solving t^2 in terms of x, we get t^2 = x-1 sub that into t^2 in the y side and simplify

OpenStudy (anonymous):

okay thank you!!

OpenStudy (gorv):

@coolkat4 t^2=x-1 t^2=y+1 implies x-1=y+1 y=x-2